OpenStudy (anonymous):

What is the center of the circle that has a diameter whose endpoints are (6, -4) and (18, 10)? is it (12, 7)?

OpenStudy (lgbasallote):

close...but no

OpenStudy (lgbasallote):

the y coordinate should be \[\huge \left(\frac{y_1 + y_2}{2} \right) \implies \left(\frac{10 - 4}{2}\right)\]

jimthompson5910 (jim_thompson5910):

somehow you did (4+10)/2 = 14/2 = 7 for the y coordinate (you lost the negative sign on the 4)

OpenStudy (anonymous):

ooh.. (12, 6)

jimthompson5910 (jim_thompson5910):

closer, but still no

OpenStudy (anonymous):

ahh.. ok i'll try again

OpenStudy (anonymous):

(12, 3)??

jimthompson5910 (jim_thompson5910):

nailed it